OpenStudy (anonymous):

What is the first step in solving 4(x + 2) = 10 + x?

OpenStudy (tommo_lcfc):

Expand the left hand side.

OpenStudy (anonymous):

Distribute by multiplying 4 by the terms in the parenthesis. Multiply by 4 on both sides. Subtract 2 on both sides. Subtract x on both sides. sorry forgot this part

OpenStudy (anonymous):

Well that pretty much explains it all =)

OpenStudy (tommo_lcfc):

Yes it does!

OpenStudy (anonymous):

i came up with x=2

OpenStudy (tommo_lcfc):

That can't be right! If we expand the brackets we get 4x+8 = 10 + x. This can be rearranged to 3x = 2 and x = 2/3, which is right.

OpenStudy (tommo_lcfc):

The previous explanation from @halo1996 explained the wrong method!

OpenStudy (anonymous):

Oh I see, I didn't read it properly - I took it to be a step-by-step, but it was asking which the first step would be. Which is distributing the 4: 4(x + 2) = 10 + x 4x + 8 = 10 + x 3x = 2 x = 2/3